# Large Language User Interfaces: Voice Interactive User Interfaces Powered by LLMs

### v1

@@ 4.3.1 Tokenization
    An input sentence is split into tokens. BERT uses WordPiece tokenization
@@ 4.3.2 Token to ID Conversion
    Each token is then mapped to an ID using BERT’s vocabulary. This can be represented as:
    $\text{ID}_{i}=\text{Vocab}(\text{"Token}_{i}\text{")},\text{ for }i=1,2,...,N$
@@
4.3.3 Segment and Positional Embeddings
@@
    For BERT, segment embeddings (to differentiate between different sentences) and positional embeddings (to give information about the position of the token in the sequence) are also added
@@
    The total embedding for a token is then:
    $E_{\text{token}}=E_{\text{word}}(ID)+E_{\text{segment}}+E_{\text{position}}$
@@
    Here, $E_{\text{word}}(ID)$ is the word embedding corresponding to the token ID, $E_{\text{segment}}$ is the segment embedding, and $E_{\text{position}}$ is the positional embedding.
    Cosine $\text{Cosine Similarity}(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$
The cosine similarity between a user’s input and an annotated node will determine the likelihood that the user desired that node. Whichever node has the highest similarity with the user’s input will end up being the classified application, and the engine will traverse its children nodes as a result. This type of Best-First Search mitigates redundant comparisons through all leaf nodes from incorrect applications, in turn significantly reducing the chance for incorrect matching with incorrect UI components.
Parameter Extraction
Proceeding down the pipeline is the core function of the engine. At this level, a powerful LLM with substantial generalization and multitasking ability is used to compare the annotated children nodes of the classified application and compare each to the user’s input. Through prompt engineering, an LLM such as Google’s T5 model fine-tuned on large text corpora for accurate NLP multitasking abilities, can determine which part of the user’s input corresponds to the current child node being checked. This extraction algorithm functions as a decision-making tool with the use of natural language processing techniques. Its objective is to analyze textual input, categorize it based on predefined criteria, and generate a JSON parameter extraction object, to be relayed to the front end and dispatched to the Redux central store.
@@
4.4 Operational Workflow
  1. 1.
    Input Analysis: The algorithm initially analyzes the input to determine its nature and relevance to each UI component child node under the classified application. This categorization is crucial in deciding the subsequent processing path
  2. 2.
    Conditional Logic: Based on the analysis, the algorithm employs conditional logic to select an appropriate extraction mechanism. Parent nodes of classified applications are checked, to guarantee the most suitable model is employed. This selection process is pivotal to the algorithm’s adaptability and accuracy in parameter extraction.
  3. 3.
    Response Generation: Depending on the categorization:
    • If the input falls into a specific category (e.g., calculations), the algorithm engages a dedicated module designed for that category, employing specialized deep-learning models, such as our custom fine-tuned Google’s T5
    • This delegation of tasks ensures we use the appropriate model for the task, employing larger more capable models for more demanding tasks such as complex arithmetic extraction, while making use of suitable smaller models for simplistic tasks. This optimization ensures inference time is not wasted on tasks that don’t require it, saving resources along with crucial response time.
  4. 4.
    JSON Output Formatting: Finally, the parameters extracted by the model are formatted into a structured JSON object, optimized and suitable for further processing or display to the end-user once sent back to the client. See Listing 3
This pipeline employed by our LLM engine allows for complex comprehension of textual input from a user, by accurately classifying and extracting parameters that the target application requires. This engine is able to streamline the connection between the user and the software by taking the intentions and requirements of both into account, utilizing its logical reasoning to complete the task as it sees fit. This is a stepping stone towards the right direction in the next generation of software.
@@ 5.1 Required Tasks
Classification is accomplished by encoding descriptive nodes for each task along with the input node using an all-MiniLM sentence-transformer model and then selecting the node with the greatest cosine similarity. This classification task is extremely reliable and the chance of failure is observed to be negligible. The task that is critical to evaluate is the parameter extraction once the input’s corresponding task has been established. Extensive experimentation and trials were conducted with plentiful transformer models from the Hugging Face hub. Models specialized in Question Answering as well as Text2Text Generation were found to perform extremely well.
@@ 5.2 Model Selection
For this evaluation, we pit the most effective models found during initial testing, with available resources for training, hosting, deployment and inference in mind)—the T5 Text2Text Generator models along with BERT and ELECTRA. The BERT and ELECTRA models were already fine-tuned on the SQuAD 2.0 dataset, which is a popular dataset used in fine-tuning models for logical reasoning and question-answering. On the other hand, the T5 model base and multi-task, trained by Google were also chosen due to their flexibility, scalability, and efficacy in abstraction and multi-task situations.
Figure 3: Accuracy comparison between models engineered for our application’s tasks. Each model was trained on logical and information extraction tasks, and further carefully prompt engineered to achieve the best performance possible for each task. Text2Text Generation Models were fine-tuned on multiple datasets to multitask excellence. The T5 custom model was specifically trained on specialized custom datasets relevant to our application’s tasks.
@@ 5.3 Observation
We can see from the following benchmarks in Figure 3 that the custom-trained T5 model does exceptionally well across all tasks, performing the best in three out of four tasks, with the last one being a close tie. The BERT model also performs very well but struggles with the logical capabilities required for the advanced calculator prompts. Google’s multitask-trained model falls extremely short in more logically involved tasks while the ELECTRA model does poorly overall. Further training of the T5 model with more diverse and better quality data would allow it to be the most versatile and powerful model for our project.
GPT Baseline VS Custom Engine Prompt Evaluation
Truth Intention GPT-4 Generated Inputs Multimodal Parser Output Eval
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
Oh, before I forget, jot down my email, it’s wanderlust.world@travelbug.com, and people shout out to me as Alex P. Globetrotter when they see me. The address? It’s 7890 Nomad Path, Backpack City. The mailman has it tough, I tell you!
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
I’m registered under the name Alex J. Turner, but everyone sends their regards to my place at 768 Rolling Rock Street, and for a quicker response, they hit me up at alex.turns@rocknmail.com.
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
Could you tell me if I need an umbrella for my walk today around the canals of Amsterdam, Netherlands?
[…]
{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
I’m packing for a business trip and I’m not sure what to bring. What’s the current weather in Cape Town, South Africa?
[…]
{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
Just finished a book set in Venice, Italy, and it got me wondering about the weather there today. Is it as sunny as in the story?
[…]
{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
My friend in Sydney mentioned they were having quite the heatwave. I’d love to compare that to the current temperatures where the Great Pyramids stand. Could you give me a quick overview of the weather there?
[…]
{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
$\times$
[…]
{"CurrentApp":"Calculator","Config":{"promptSequence": "24 / 6"}}
I’ve got 24 cupcakes, and I need to divide them evenly among my 6 friends. How many does each person get?
[…]
{"CurrentApp":"Calculator","Config":{"promptSequence": "24/6"}}
$\checkmark$
Table 2: Comparison of GPT-4 generated inputs with our custom multimodal LLM engine consisting of custom fine-tuned Google T5 & BERT-Large-Uncased-Whole-Word-Masking-Squad2 outputs. The model performs with extreme efficacy on descriptive and lengthy prompts. However, it can be observed to struggle with advanced mathematical prompts as it is strictly limited to segment extraction. Mathematical prompts can come in all sorts of wrangled manners, and the BERT model cannot handle expression extraction with logical implications. This is why the custom-tuned T5 proves to be so powerful, with all tasks.
@@ 6.0.2 LLM Engine Embellishments
Exploring configurations regarding transformer models that classify, extract and parse information in the LLM engine will improve accuracy and inference speed. The choice between delegating tasks to different models versus using a singular multi-task model for all tasks is open to be explored.
@@ 6.0.3 AutoGen
Expanding on the back-end embellishments, making use of Microsoft’s AutoGen (Wu et al. (2023)), Mistral’s Mixtral 8x7B mixture of experts model (Jiang et al. (2024)), or Ollama’s model hot-swapping to allow for multi-model agent setups may be the ultimate solution for delegating a task to the appropriate model rather than having a singular model handle all tasks. If an AutoGen setup can delegate a task to the best-suited LLM available in the engine, accuracy can be maximized across all tasks and inference times can be dynamic depending on the complexity of the task, rather than potentially sluggish times even for simpler tasks that don’t require larger models.
@@ 6.0.4 Concurrent Processing
Further scaling would call for large-scale concurrent processing of incoming user requests. Whilst our backend server is currently multi-threaded and can handle many inputs simultaneously, further scaling to incorporate asynchronous processing with expanded capabilities, along with more robust load balancing and distributed computing systems, would bulletproof our framework’s servers.

### v2

@@ 4.3 Parsing & Mapping User Input
    An input sentence is split into tokens, adhering to an agent’s specific LLM encoding requirements. We will cover BERT and its use of WordPiece tokenization for this example
@@
    Each token is then mapped to an ID using BERT’s vocabulary. This can be represented as
    $\text{ID}_{i}=\text{Vocab}(\text{``Token}_{i}\text{")},\text{ for }i=1,2,...,N$
@@
  2. 2.
    The total embedding for a token, $E_{token}$, includes word, segment, and positional embeddings.
4.3.1 Mapping via Classification
To classify user input, we encode the input and node representations:
  • For user input $u_{i}:v_{u_{i}}=encode(tokenize(u_{i}))$
  • For a node $n_{j}$ representing an application or component: $v_{n_{j}}=encode(tokenize(n_{j}))$
Let U denote the user’s encoded input:
$U=\{u_{1},u_{2},u_{3}...\}$
The system then maps user input to system nodes using an interpretation function that performs classification, $f:U\longrightarrow S$, where:
$f(U_{i})=n_{k}$, where $n_{k}$ is chosen such that:
$n_{k}=argmax(cosine\_similarity(\vec{v}_{u_{i}},\vec{v}_{n_{j}}))$
This process classifies the user’s input to the most relevant application component in the annotation tree. To handle complex inputs that involve specific actions or components, we extend the function $f$ to further map inputs to nested sets (components) within A, W, or C:
$f(u_{i})=a_{j}\in A\textit{ or }f(u_{i})=w_{j}\in W\textit{ or }f(u_{i})=c_{j}\in C$
This implies that the function not only identifies the relevant application but also the specific action or component within that application.
4.3.2 Entity Extraction and Parameterization via LLM Inference
At this level, the custom fine-tuned LLM agents with specialized ability such as Google’s T5 model trained on specific tasks from our library in a teacher-student environment, enable accurate parsing of complex user inputs into structured data that the system can act upon. This process is crucial for interpreting the specific details contained within the user’s input and mapping them to actionable UI component parameters in our system.
Figure 5: Entity Extraction & Paramaterization
For inputs that require extracting parameters or entities, we define a transformation $g$: $U$$P$, where P is a set of parameters extracted from the inputs. This can be used alongside $f$ to not only map an input to an application component but also extract the necessary information to perform the action:
$\textit{g}(\textit{u}_{\textit{i}})=\{\textit{p}1,\textit{p}2,…\}$
where $u_{i}$ is the encoded user input, and $\{p_{1},p_{2},...\}$ are the extracted parameters. These parameters are then used to dynamically adapt the system’s response to the user’s intent. This extraction algorithm functions as a decision-making tool with the use of natural language processing techniques. Its objective is to analyze textual input, categorize it based on predefined criteria, and generate a JSON parameter extraction object to be relayed to the front end and dispatched to the Redux central store.
4.3.3 Action Execution Based on Mapping
Finally, we use the outcomes of f and $g$ to create a final JSON object and execute actions within the system. This involves sending the final JSON object to the front end where the Redux central store invokes specific methods associated with the identified components, passing the extracted parameters and dynamically updating the UI. Let the following represent the introduction of relevant methods:
$execute(f(u_{i}),g(u_{i}))$
4.3.4 Example:
Assume a user input $u_{i}$= “I have a flight to France, what’s it like out there?". The interpretation function f maps this input to the weather application’s “location" field component $w_{location}\in W$, and $g$ extracts “France" as a parameter. The system then executes the location-based action for “France" based on this mapping and parameter extraction.
The defined detailed approach allows the system to dynamically interpret user inputs by understanding their intent and mapping them to the most relevant application and components in the application library system. It ensures that user inputs are handled efficiently, navigating through the hierarchical structure of the application library based on semantic similarity, thereby improving user experience through accurate identification and interpretation of a user’s needs.
4.4 Operational Workflow
@@
    Input Analysis The algorithm initially analyzes the input to determine its nature and relevance to each UI component child node under the classified application. This categorization is crucial in deciding the subsequent processing path.
    1. (a)
      Top-Down, First-Match Traversal The system starts the mapping at the root of the application tree. It uses our defined interpretation function f to select the child node with the highest cosine similarity to the user input. The process repeats until a leaf node is reached, moving to children nodes only after the parent node is classified.
    2. (b)
      Early Stopping for Efficiency & Error Mitigation A threshold is in place for cosine similarity. If the similarity between the user input and the best-matching node at any level is below a predefined threshold, the search is halted, and the system can proceed with action execution based on the current best match, or prompt the user to try again. This optimizes the search process, ensures responsiveness and mitigates errors that can result from forced mapping.
@@
    Conditional Logic Based on the analysis, the algorithm employs conditional logic to deploy appropriate agents in the extraction mechanism. Parent nodes of classified applications are checked, to guarantee the most suitable model is utilized. This selection process is pivotal to the algorithm’s adaptability and accuracy in parameter extraction.
@@
    Response Generation Depending on the categorization:
    • If the input falls into a specific category (e.g. calculations), the algorithm engages a dedicated module designed for that category, employing specialized deep-learning models, such as our custom fine-tuned Google’s T5
    • This delegation of tasks makes it so that we use the most appropriate model for the task, employing larger more capable models for more demanding tasks such as complex arithmetic extraction, while making use of suitable smaller models for simplistic tasks. This optimization ensures valuable inference time from larger models is not wasted on tasks that don’t require it, saving resources along with crucial response time.
  4. 4.
    JSON Output Formatting Finally, the parameters extracted by the model are formatted into a structured JSON object, optimized and suitable for further processing or for display to the end-user once sent back to the client. See Listing 1.3.
@@
This pipeline employed by our LLM engine allows for complex comprehension of textual input from a user, by accurately classifying and extracting parameters that the target application requires. Our engine is able to streamline the connection between the user and the software by accounting for the intentions and requirements of both, utilizing its logical reasoning to complete the task as it sees fit. This is a stepping stone in the right direction for the next generation of software.
@@ 5 Evaluation
{adjustwidth}
[…]
Table 2: Comparison of GPT-4 generated inputs with our custom multimodal LLM engine consisting of custom fine-tuned Google T5 & BERT-Large-Uncased-Whole-Word-Masking-Squad2 outputs. The model performs with extreme efficacy on descriptive and lengthy prompts.
@@ 5.1 Required Tasks
Classification is accomplished by encoding descriptive nodes for each task along with the input node using an all-MiniLM sentence-transformer model and then selecting the node with the greatest cosine similarity. This classification task is extremely reliable and the chance of failure is observed to be negligible. The critical task to evaluate is the parameter extraction once the input’s corresponding application has been established. Extensive experimentation and trials were conducted with plentiful transformer models from the Hugging Face hub. Models specialized in Question Answering as well as Text2Text Generation were found to perform extremely well.
@@ 5.2 Model Selection
For this evaluation, we pit the most effective models found during initial testing, with available resources for training, hosting, deployment and inference in mind—the T5 Text2Text Generator models along with BERT and ELECTRA. The BERT and ELECTRA models were already fine-tuned on the SQuAD 2.0 dataset, which is a popular dataset used in fine-tuning models for logical reasoning and question-answering. On the other hand, the T5 model base and multi-task, trained by Google, were also chosen due to their flexibility, scalability, and efficacy in abstraction and multi-task situations.
Figure 6: Accuracy comparison between models engineered for our application’s tasks. Each model was trained on logical and information extraction tasks, and further carefully prompt engineered to achieve the best performance possible for each task.
@@ 5.3 Observation
We can see from the benchmarks in Figure 6 that the T5 model trained on specialized custom datasets relevant to specific application tasks does exceptionally well across all tasks. It performs the best in three out of four tasks, with the last one being a close tie. The BERT model also performs very well but struggles with the logical capabilities required for the advanced calculator prompts. Google’s multitask-trained model falls extremely short in more logically involved tasks, while the ELECTRA model does poorly overall. Further training of the T5 model with more diverse and better quality data would allow it to be the most versatile and powerful model for our framework.
@@ 6.0.2 LLM Engine Embellishments
Exploring configurations regarding transformer models that classify, extract, and parse information in the LLM engine will improve accuracy and inference speed. The choice between delegating tasks to different models versus using a singular multi-task model for all tasks is open to be explored.
@@ 6.0.3 AutoGen
Expanding on the back end embellishments, making use of Microsoft’s AutoGen [11], Mistral’s Mixtral 8x7B mixture of experts model [12], or Ollama’s model hot-swapping to allow for multi-model agent setups may be the ultimate solution for delegating a task to the appropriate model. AutoGen’s optimized mixture of experts system would pose many benefits over a manual swapping system, especially with the capability for agents to communicate with one another. If an AutoGen setup can delegate a task to the best-suited LLM amongst each other, accuracy can be maximized across all tasks and inference times can be further optimized.
@@ 6.0.4 Concurrent Processing
Further scaling would call for large-scale concurrent processing of incoming user requests. Whilst our back-end server is currently multi-threaded and can handle many inputs simultaneously, further scaling to incorporate asynchronous processing with expanded capabilities, along with more robust load balancing and distributed computing systems, would bulletproof our framework’s servers.
7 Conclusion
Our proposed LM-UI framework re-imagines the development process of contemporary user interface systems. We employ a system that automates the creation of textual semantic UI component annotations, which are supplementary to an agent-based prompting back end. This hybrid approach harnesses the power of LLMs to map generic, vague, and convoluted user textual input to their relevant components. We can further parse the input through logical entity extraction processes, creating a JSON object capable of dynamically adapting the UI. Our evaluations showcase the accuracy and precision of this approach, allowing us to take a large step closer to bringing what was once a science fiction dream, to reality.
Furthermore, the proposed framework is just the beginning. The extensions and features we can implement atop this core framework increase the benefits substantially. From streamlining complex workflows in industries such as finance and healthcare to revolutionizing highly productive multitasking environments in modern workplaces, the implications are vast and worthy of exploration. We aim to continue our research efforts, to inspire and propagate a new wave of research and development in this direction. Our objective is to fully leverage the potential of LLMs in creating more adaptive, intuitive, and efficient software systems.